# Capturing Human Motion from Monocular Images in World Space with Weak-supervised Calibration

### v4

@@ Abstract
Previous methods for 3D human motion recovery from monocular images often fall short due to reliance on camera coordinates, leading to inaccuracies in real-world applications where complex shooting conditions are prevalent. The limited availability and diversity of focal length labels further exacerbate misalignment issues in reconstructed 3D human bodies. To address these challenges, we introduce W-HMR, a weak-supervised calibration method that predicts “reasonable” focal lengths based on body distortion information, eliminating the need for precise focal length labels. Our approach enhances 2D supervision precision and recovery accuracy. Additionally, we present the OrientCorrect module, which corrects body orientation for plausible reconstructions in world space, avoiding the error accumulation associated with inaccurate camera rotation predictions. Our contributions include a novel weak-supervised camera calibration technique, an effective orientation correction module, and a decoupling strategy that significantly improves the generalizability and accuracy of human motion recovery in both camera and world coordinates. The robustness of W-HMR is validated through extensive experiments on various datasets, showcasing its superiority over existing methods. Codes and demos have been made available on the project page https://yw0208.github.io/w-hmr/.
Index Terms:
Human motion recovery, monocular 3D reconstruction, world space, camera calibration, motion capture
@@
Figure 2: Comparison among the camera models used by the other four models [1, 2, 3, 4] and our W-HMR. $f_{crop}$ and $f_{full}$ denote the focal lengths of the full and cropped images, respectively. $R_{c}$ is the camera rotation matrix, and $I$ is the identity matrix. $s$ means scale. $h_{bbox}$ and $h$ are the heights of the bounding box. But $h_{bbox}$ is dynamic, and $h$ is a constant value. $t_{b}^{z}$ refers to the vertical distance from the human body to the plane where the camera is located. The red items mean trainable, i.e., predicted by neural networks.
I Introduction
Recovering 3D human motion from monocular images is crucial in human-centered research, with wide-ranging applications in try-on, animation, and virtual reality. However, there is a noticeable lack of research on global body reconstruction from images captured under complex shooting conditions. For previous methods based on the camera coordinate, the reconstructed bodies outside controlled environments are often unsuitable for real-world applications, as depicted in Fig. 1.
The primary challenge in this task arises from the limited camera information. Among camera parameters, focal length is critical in distortion effects, influencing the alignment of reconstructed 3D human bodies with original images. There is nothing inherently wrong with this misalignment. However, since most existing models are trained with 2D joint labels, this misalignment due to distortion affects the training and ultimately leads to decreased recovery accuracy. Additionally, camera rotation determines the reasonableness of the reconstructed body pose in the world (global) or camera (local) coordinate after coordinate transformation. Fig. 2 illustrates the treatment of the camera in existing models, emphasizing the need for improved methods. The focal length either involves simple manual definitions as seen in HMR [1] and CLIFF [4], or network predictions as utilized by SPEC [3] and Zolly [2].
As for camera rotation, most models [5, 6, 7, 8, 9, 10, 11, 12] do not consider it and recover body in the camera coordinate, except for SPEC. Even based on the most comprehensive camera model, SPEC still has limitations due to the inherent flaws in its CamCalib model, which are discussed below.
Firstly, according to traditional methods like SPEC, the focal length is predicted based on background environmental information, which cannot handle finer distortions of body parts and is inadequate for the body recovery task. Moreover, focal length labels are always rare and not diverse enough, which can not fit the need for model training. To address these challenges, we propose a weak-supervised camera calibration method. The core idea is that we do not need to predict original focal lengths but the “reasonable” ones. Our method leverages body distortion information to predict a reasonable focal length for perspective projection. It eliminates the need for supervision with rare focal length labels and makes it more suitable for body tasks. Based on the full-perspective camera model, 2D supervision becomes more precise, improving recovery accuracy.
Secondly, the predicted camera rotation is not entirely accurate. With pseudo camera rotation as input, training a model like SPEC to predict poses in the world coordinate would lead to error accumulation. The error in pseudo camera parameters would accumulate during training and impact the final recovery precision. Consequently, the accuracy of recovery is limited. Additionally, it decreases the model’s generalizability. For example, compared to SPEC’s advantages in world space, it performs poorly in the camera coordinate, as shown in Tab. II. We adopt an indirect approach for predicting body poses in the world coordinate. First, we train our model to achieve optimal pose and shape accuracy in the camera coordinate. Then, we propose a novel module named OrientCorrect to assist in correcting body orientation, thus avoiding the issues above and enabling human motion recovery in the world coordinate.
@@ I Introduction
    We present the first weak-supervised camera calibration method focus on the human recovery task. We achieve reasonable focal length prediction without focal length labels but rely on information about the human body distortion. Better mesh-image alignment makes 2D supervision more precise and effectively improves recovery accuracy. Our method even outperforms the latest supervised method like Zolly [2].
@@
II Related Work
II-A Camera Model in Human Recovery
To visualize the reconstruction results and utilize 2D labels, 3D human recovery models require a camera model. The classical HMR [1] employed a weak-perspective camera model, which assumes a considerable focal length of 5000 pixels to mitigate human distortion. This assumption is suitable for images captured with a long focal length when the human subject is distant from the camera. Due to its simplicity, the weak-perspective camera model has been widely adopted in subsequent works [13, 7, 11, 14, 10].
To our knowledge, Kissos et al. [15] was the first to enhance the weak-perspective camera model in pose estimation. They assumed a constant camera Field of View (FoV) of 55 degrees instead of a fixed focal length. Later, CLIFF [4] also adopted the same camera model. SPEC [3] took a different approach by constructing a new dataset and training a camera parameter prediction model named CamCalib. The latest method, Zolly [2], indirectly predicts the focal length through a well-designed translation prediction module, albeit partially relying on focal length labels and only predicting the focal length of cropped images.
@@ II-A Camera Model in Human Recovery
II-B Regression-based Method
@@
In the 3D human body reconstruction domain, methods can be broadly classified into regression-based and optimization-based approaches. Optimization-based methods are beyond the scope of this paper and can be explored in the literature review [16]. Regression-based models can be further categorized into model-based and model-free approaches based on the output type. Model-based models refer to those that output template parameters, with the majority of current models [1, 14, 13, 10, 11, 17] still relying on the SMPL algorithm [18], and a few [19, 20, 21, 22] based on SCAPE [23], GHUM&GHUML [24], and SMPL-X [25]. These models leveraged the prior knowledge embedded in parameter models but were limited to reconstructing finer motion details due to constraints in degrees of freedom.
@@
In contrast, model-free models in this field output voxels [26, 27], UV position maps [28, 29] and 3D vertex coordinates [12, 30, 31, 32]. Here, we mainly focus on methods that output 3D vertex coordinates. Apart from methods that output voxels, all other types of output would be converted to 3D vertices and then transformed into meshes based on predefined triangle faces. One classic model-free method, Graphormer [12], designed a regression module based on transformer [33], which directly outputs 431 3D sparse vertices. The advantage of model-free models is that they can recover more details than model-based models. However, their limitations lie in their inability to effectively utilize prior knowledge of the human body, and their models tend to have larger sizes. To overcome the limitations of both model-based and model-free approaches, we propose a hybrid method where the model-based regression module provides a strong initialization, and the model-free module further refines the results. This approach aims to achieve optimal mesh precision with relatively fewer parameters, bridging the gap between model-based and model-free methods.
@@
Figure 3: Pipeline of W-HMR. A PyMAF-like backbone is adopted to extract features and predict the recovery results in the camera coordinate. ”glob info” refers to global information consisting of five elements. $(\varDelta c_{x},\,\varDelta c_{y})$ are offsets of the bounding box center relative to the center of the original image. FULL2 2D supervision refers to 2D joint supervision on full images, which is based on full-perspective camera model. The CamClib takes the whole image as input and outputs camera rotation matrices. The OrientCorrect takes three vectors as input and outputs body orientation in the world coordinate. $\{s,\,t_{x},\,t_{y}\}$ are scales and translations in the camera coordinate. $f$ is focal length and $(t_{b}^{x},\,t_{b}^{y},\,t_{b}^{z})$ is the translation to camera in the world coordinate. Note $\varPhi_{3}$ and $\phi_{3}^{s}$ are features extracted by the PyMAF-like backbone. If you are confused about our feature extraction, please refer to Sec. IV-A for our network details.
III Method
@@
The overview of W-HMR is shown in Fig. 3. The input image undergoes a PyMAF-like model to extract features and obtain the recovered human motion in the camera coordinates. We design a focal length predictor to predict the focal length to implement a full-perspective projection on the full image for 2D joint supervision. In addition, we propose the OrientCorrect module to correct the body orientation to keep the pose reasonable in the world space. Our method is introduced in five sections. We first present some prior knowledge in Sec. III-A. This explains why we need to use the full perspective camera model and the full image information. In Sec. III-B, we focus on how we predict the focal length and how we use the predicted focal length for supervision. In Sec. III-C, we describe how OrientCorrect works, another essential module. In Sec. III-D, we describe how to stabilize model training without focal length labels. Finally, in Sec. III-E, we add some introduction to more loss functions. To fully demonstrate the effectiveness of our method, we experiment with various network structures. Specific details can be found in Sec. IV-A.
III-A Preliminary
@@
Translation and Scale Following  [1], we continue to utilize SMPL parameters $\varTheta$ as our template. The SMPL parameters encompass pose $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$, shape $\boldsymbol{\beta}\in\mathbb{R}^{10}$, where $\boldsymbol{\theta}_{0}$ is the body orientation. Specifically, pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$ can be inputted into a pre-defined function $M\left(\boldsymbol{\theta},\boldsymbol{\beta}\right)\in\mathbb{R}^{N\times 3}$ to obtain $N$ vertex coordinates, i.e., $(X,Y,Z)$ we talked about before. The mesh derived from these vertices is the reconstructed human mesh. To project it onto the image, the model outputs three additional parameters, i.e., translation $(t_{x},t_{y})$, and scale $\boldsymbol{s}$. As for $\boldsymbol{s}$ and $(t_{x},t_{y})$, they determine the size and position of the human meshes within the cropped image. Based on the usual assumptions, we assume a human body range of 2 meters, which establishes the relationship between $t_{b}^{z}$ and $f$ as
@@
where $h_{bbox}$ refers to the height of the bounding box. This formula means that the 2-meter body range is imaged on the image, and the range becomes scale $s$ multiplied by the height of the bounding box $h_{bbox}$. Scale is why this formula is general. The scale is predicted by a neural network, which represents the size of the human in the bounding box. Therefore, it is dynamically adaptive, making this formula valid in most cases. The evidence is qualitative results in our paper. Our recovered mesh can be reprojected to overlap well with the human area. In addition, much work has adopted this assumption and this formula since it was proposed in 2018 [1]. For a more detailed reasoning process, please refer to [4] or [15]. Our goal is to leverage the relationship between $t_{b}^{z}$ and $f$ to indirectly predict $f$ by getting $t_{b}^{z}$, thus achieving full perspective.
III-B Weak-Supervised Camera Calibration
Focal length prediction is essentially an ill-posed problem. From Sec. III-A, we can see too many influencing factors in the process, from 3D objects to 2D images. Predicting the actual focal length through a neural network is almost impossible. So, we propose using human body distortion information to predict a ”reasonable” focal length for full-perspective projection.
Camera Calibration The focal length predictor takes the final feature map $\varPhi_{3}$ as input and output a single value $t_{b}^{z}$. Please refer to Sec.  IV-A for how to extract features $\varPhi_{3}$. To constrain the range of the predicted $t_{b}^{z}$ [2], we employ a sigmoid function and scale it by 10. This adjustment is necessary because distortion becomes irrelevant when a person stands 10 meters from the camera. We add batch normalization to avoid gradient vanishing or exploding caused by the sigmoid function. The focal length can be calculated according to $f=\frac{sh_{bbox}t_{b}^{z}}{2}$, where the height of bounding box $h_{bbox}$ is known and scale $s$ is output by the PyMAF-like regression module. From Equ. 2, to implement the projection, we also need to know the translation of the body to the camera. $t_{b}^{z}$ has already been generated, so we need further to compute $(t_{b}^{x},t_{b}^{y})$. Following [15], $(t_{b}^{x},t_{b}^{y})=(t_{x},t_{y})$ when projecting in cropped image. But in the full image, the conversion functions for $(t_{x},t_{y})$ to $(t_{x}^{b},t_{y}^{b})$ are summarized below

### v6

@@ Abstract
Previous methods for 3D human motion recovery from monocular images often fall short due to reliance on camera coordinates, leading to inaccuracies in real-world applications. The limited availability and diversity of focal length labels further exacerbate misalignment issues in reconstructed 3D human bodies. To address these challenges, we introduce W-HMR, a weak-supervised calibration method that predicts “reasonable” focal lengths based on body distortion information, eliminating the need for precise focal length labels. Our approach enhances 2D supervision precision and recovery accuracy. Additionally, we present the OrientCorrect module, which corrects body orientation for plausible reconstructions in world space, avoiding the error accumulation associated with inaccurate camera rotation predictions. Our contributions include a novel weak-supervised camera calibration technique, an effective orientation correction module, and a decoupling strategy that significantly improves the generalizability and accuracy of human motion recovery in both camera and world coordinates. The robustness of W-HMR is validated through extensive experiments on various datasets, showcasing its superiority over existing methods. Codes and demos have been made available on the project page https://yw0208.github.io/w-hmr/.
Keywords:
Human motion recovery Monocular 3D reconstruction World space Camera calibration Motion capture
1 Introduction
Recovering 3D human motion from monocular images is crucial in human-centered research, with wide-ranging applications in try-on, animation, and virtual reality. However, there is a noticeable lack of research on global body reconstruction from images captured under complex shooting conditions. Previous methods based on the camera coordinate can only be used in a controlled laboratory environment. When faced with images taken at various strange angles in the wild, previous methods often cannot recover useful results, which can be seen in Fig. 1.
@@ 1 Introduction
The primary challenge in this task arises from the limited camera information. In the camera parameters, there is an intrinsic parameter focal length that cannot be decoupled. This unknown parameter causes misalignment of the recovered human mesh and the original images. Since most existing models are trained with 2D joint labels, this misalignment due to the unknown focal lengths affects the training and ultimately leads to decreased recovery accuracy. As for extrinsic parameters, camera rotations are essential for coordinate transformation. When the shooting angle makes the world coordinates and camera coordinates inconsistent, the unknown camera rotation limits the application of previous methods based on the camera coordinate.
@@
Figure 2: Comparison among the camera models used by the other four models (Kanazawa et al., 2018; Wang et al., 2023; Kocabas et al., 2021b; Li et al., 2022b) and our W-HMR. $f_{crop}$ and $f_{full}$ denote the focal lengths of the full and cropped images, respectively. $R_{c}$ is the camera rotation matrix, and $I$ is the identity matrix. $s$ means scale. $h_{bbox}$ and $h$ are the heights of the bounding box. But $h_{bbox}$ is dynamic, and $h$ is a constant value. $t_{b}^{z}$ refers to the vertical distance from the human body to the plane where the camera is located. The red items mean trainable, i.e., predicted by neural networks.
Fig. 2 illustrates the treatment of the camera in existing models, emphasizing the need for improved methods. The focal length either involves simple manual definitions as seen in HMR (Kanazawa et al., 2018) and CLIFF (Li et al., 2022b), or network predictions as utilized by SPEC (Kocabas et al., 2021b) and Zolly (Wang et al., 2023). As for camera rotation, most models (Zhou et al., 2010; Guan et al., 2009; Zhang et al., 2023a, 2020a; Kolotouros et al., 2019a; Kocabas et al., 2020, 2021a; Lin et al., 2021b) do not consider it and recover body in the camera coordinate, except for SPEC. SPEC, including recent work (Shin et al., 2024; Zhang et al., 2023b), uses a neural network to predict camera rotation to assist human mesh recovery in the world coordinate. Even based on the most comprehensive camera model, SPEC still has limitations due to its inherent flaws. Using predicted camera rotation as input to inference for world coordinate pose prediction leads to error accumulation, limiting recovery accuracy and reducing model generalizability. A piece of evidence is that SPEC is poor in the camera coordinate compared to its advantage in world space (see Tab. 1). To address the above defects, we propose the following two technical contributions.
First, traditional methods such as SPEC predict focal length based on background environment information. This kind of method can predict a coarse focal length to alleviate the problem of misalignment. However, it cannot obtain subtle distortions of body parts and cannot meet the needs of human-centric tasks. In addition, focal length labels are difficult to obtain and not diverse enough to meet the needs of model training. In view of the coarseness of predicted focal lengths and the shortage of labels, we propose a weak-supervised camera calibration method. Specifically, we use more refined mesh-image alignment features to predict focal length. In addition, we use the projection relationship between 3D and 2D joints to supervise the training of our focal length predictor. We let the focal length predictor learn how to predict a suitable focal length to project the predicted 3D joints onto the image and match the 2D joint positions. Compared with traditional methods, our method can capture more detailed misalignment information and is more suitable for human-centric tasks. With our method, 2D supervision becomes more precise and the recovery accuracy is improved.
Second, we adopt a two-stage decoupling method to predict body pose in the world coordinate. We decouple the camera parameters into intrinsics and extrinsics, and the human pose into global body orientation and local joint rotations. In the first stage, we only consider the intrinsic parameter and joint rotations to achieve the best recovery performance in the camera coordinate. Then, we propose a new OrientCorrect module that uses pseudo camera rotation, global information, local joint rotations, and human features to predict global body orientation. Compared with previous methods, the independent OrientCorrect does not harm the existing accuracy of the model. Our method mitigates the impact of incorrect camera rotation on recovery accuracy and achieves higher accuracy in different coordinates simultaneously.
@@
    We present the first weak-supervised camera calibration method focus on the human recovery task. We achieve reasonable focal length prediction without focal length labels but rely on information about the human body distortion. Better mesh-image alignment makes 2D supervision more precise and effectively improves recovery accuracy. Our method even outperforms the latest supervised method like Zolly (Wang et al., 2023).
@@
2 Related Work
2.1 Camera Model in Human Recovery
To visualize the reconstruction results and utilize 2D labels, 3D human recovery models require a camera model. The classical HMR (Kanazawa et al., 2018) employed a weak-perspective camera model, which assumes a considerable focal length of 5000 pixels to mitigate human distortion. This assumption is suitable for images captured with a long focal length when the human subject is distant from the camera. Due to its simplicity, the weak-perspective camera model has been widely adopted in subsequent works (Zhang et al., 2021a, 2023a; Kocabas et al., 2021a; Goel et al., 2023; Kocabas et al., 2020).
To our knowledge, Kissos et al. (Kissos et al., 2020) was the first to enhance the weak-perspective camera model in pose estimation. They assumed a constant camera Field of View (FoV) of 55 degrees instead of a fixed focal length. Later, CLIFF (Li et al., 2022b) also adopted the same camera model. SPEC (Kocabas et al., 2021b) took a different approach by constructing a new dataset and training a camera parameter prediction model named CamCalib. The latest method, Zolly (Wang et al., 2023), indirectly predicts the focal length through a well-designed translation prediction module, albeit partially relying on focal length labels and only predicting the focal length of cropped images.
@@ 2.1 Camera Model in Human Recovery
2.2 Regression-based Method
@@
In the 3D human body reconstruction domain, methods can be broadly classified into regression-based and optimization-based approaches. Optimization-based methods are beyond the scope of this paper and can be explored in the literature review (Tian et al., 2023). Regression-based models can be further categorized into model-based and model-free approaches based on the output type. Model-based models refer to those that output template parameters, with the majority of current models (Kanazawa et al., 2018; Goel et al., 2023; Zhang et al., 2021a; Kocabas et al., 2020, 2021a; Li et al., 2023) still relying on the SMPL algorithm (Loper et al., 2015), and a few (Pons-Moll et al., 2015; Zuffi and Black, 2015; Zanfir et al., 2021a, b) based on SCAPE (Anguelov et al., 2005), GHUM&GHUML (Xu et al., 2020), and SMPL-X (Pavlakos et al., 2019). These models leveraged the prior knowledge embedded in parameter models but were limited to reconstructing finer motion details due to constraints in degrees of freedom.
@@
In contrast, model-free models in this field output voxels (Varol et al., 2018; Zheng et al., 2019), UV position maps (Zhang et al., 2020b; Shetty et al., 2023) and 3D vertex coordinates (Lin et al., 2021b; Moon and Lee, 2020; Lin et al., 2021a; Choi et al., 2020). Here, we mainly focus on methods that output 3D vertex coordinates. Apart from methods that output voxels, all other types of output would be converted to 3D vertices and then transformed into meshes based on predefined triangle faces. One classic model-free method, Graphormer (Lin et al., 2021b), designed a regression module based on transformer (Dosovitskiy et al., 2020), which directly outputs 431 3D sparse vertices. The advantage of model-free models is that they can recover more details than model-based models. However, their limitations lie in their inability to effectively utilize prior knowledge of the human body, and their models tend to have larger sizes. To overcome the limitations of both model-based and model-free approaches, we propose a hybrid method where the model-based regression module provides a strong initialization, and the model-free module further refines the results. This approach aims to achieve optimal mesh precision with relatively fewer parameters, bridging the gap between model-based and model-free methods.
@@
Figure 3: Pipeline of W-HMR. A PyMAF-like backbone is adopted to extract features and predict the recovery results in the camera coordinate. “glob info” refers to global information consisting of five elements. $(\varDelta c_{x},\,\varDelta c_{y})$ are offsets of the bounding box center relative to the center of the original image. FULL2 2D supervision refers to 2D joint supervision on full images, which is based on full-perspective camera model. The CamClib takes the whole image as input and outputs camera rotation matrices. The OrientCorrect takes three vectors as input and outputs body orientation in the world coordinate. $\{s,\,t_{x},\,t_{y}\}$ are scales and translations in the camera coordinate. $f$ is focal length and $(t_{b}^{x},\,t_{b}^{y},\,t_{b}^{z})$ is the translation to camera in the world coordinate. Note $\varPhi_{3}$ and $\phi_{3}^{s}$ are features extracted by the PyMAF-like backbone. If you are confused about our feature extraction, please refer to Sec. 4.1 for our network details.
3 Method
@@
The overview of W-HMR is shown in Fig. 3. The input image undergoes a PyMAF-like model to extract features and obtain the recovered human motion in the camera coordinates. We design a focal length predictor to predict the focal length to implement a full-perspective projection on the full image for 2D joint supervision. In addition, we propose the OrientCorrect module to correct the body orientation to keep the pose reasonable in the world space. Our method is introduced in five sections. We first present some prior knowledge in Sec. 3.1. This explains why we need to use the full perspective camera model and the full image information. In Sec. 3.2, we focus on how we predict the focal length and how we use the predicted focal length for supervision. In Sec. 3.3, we describe how OrientCorrect works, another essential module. In Sec. 3.4, we describe how to stabilize model training without focal length labels. Finally, in Sec. 3.5, we add some introduction to more loss functions. To fully demonstrate the effectiveness of our method, we experiment with various network structures. Specific details can be found in Sec. 4.1.
3.1 Preliminary
@@
Translation and Scale Following  (Kanazawa et al., 2018), we continue to utilize SMPL parameters $\varTheta$ as our template. The SMPL parameters encompass pose $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$, shape $\boldsymbol{\beta}\in\mathbb{R}^{10}$, where $\boldsymbol{\theta}_{0}$ is the body orientation. Specifically, pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$ can be inputted into a pre-defined function $M\left(\boldsymbol{\theta},\boldsymbol{\beta}\right)\in\mathbb{R}^{N\times 3}$ to obtain $N$ vertex coordinates, i.e., $(X,Y,Z)$ we talked about before. The mesh derived from these vertices is the reconstructed human mesh. To project it onto the image, the model outputs three additional parameters, i.e., translation $(t_{x},t_{y})$, and scale $\boldsymbol{s}$. As for $\boldsymbol{s}$ and $(t_{x},t_{y})$, they determine the size and position of the human meshes within the cropped image. Based on the usual assumptions, we assume a human body range of 2 meters, which establishes the relationship between $t_{b}^{z}$ and $f$ as
@@
where $h_{bbox}$ refers to the height of the bounding box. This formula means that the 2-meter body range is imaged on the image, and the range becomes scale $s$ multiplied by the height of the bounding box $h_{bbox}$. Scale is why this formula is general. The scale is predicted by a neural network, which represents the size of the human in the bounding box. Therefore, it is dynamically adaptive, making this formula valid in most cases. The evidence is qualitative results in our paper. Our recovered mesh can be reprojected to overlap well with the human area. In addition, much work has adopted this assumption and this formula since it was proposed in 2018 (Kanazawa et al., 2018). For a more detailed reasoning process, please refer to (Li et al., 2022b) or (Kissos et al., 2020). Our goal is to leverage the relationship between $t_{b}^{z}$ and $f$ to indirectly predict $f$ by getting $t_{b}^{z}$, thus achieving full perspective.
3.2 Weak-Supervised Camera Calibration
Focal length prediction is essentially an ill-posed problem. From Sec. 3.1, we can see too many influencing factors in the process, from 3D objects to 2D images. Predicting the actual focal length through a neural network is almost impossible. So, we propose using human body distortion information to predict a “reasonable” focal length for full-perspective projection.
Camera Calibration The focal length predictor takes the final feature map $\varPhi_{3}$ as input and output a single value $t_{b}^{z}$. Please refer to appendices for how to extract features $\varPhi_{3}$. To constrain the range of the predicted $t_{b}^{z}$ (Wang et al., 2023), we employ a sigmoid function and scale it by 10. This adjustment is necessary because distortion becomes irrelevant when a person stands 10 meters from the camera. We add batch normalization to avoid gradient vanishing or exploding caused by the sigmoid function. The focal length can be calculated according to $f=\frac{sh_{bbox}t_{b}^{z}}{2}$, where the height of bounding box $h_{bbox}$ is known and scale $s$ is output by the PyMAF-like regression module. From Equ. 2, to implement the projection, we also need to know the translation of the body to the camera. $t_{b}^{z}$ has already been generated, so we need further to compute $(t_{b}^{x},t_{b}^{y})$. Following (Kissos et al., 2020), $(t_{b}^{x},t_{b}^{y})=(t_{x},t_{y})$ when projecting in cropped image. But in the full image, the conversion functions for $(t_{x},t_{y})$ to $(t_{x}^{b},t_{y}^{b})$ are summarized below